OpenStudy (anonymous):

Let u = <-3, 4>, v = <8, 2>. Find u + v

OpenStudy (amistre64):

add like parts maybe?

OpenStudy (anonymous):

That's what I was thinking. so <-3+8 ; 4+2> <5, 6> ?

OpenStudy (amistre64):

perfect .... thought i typed that

OpenStudy (anonymous):

I believe that is the answer. Let u = <3, -1>, v = <-6, -6>. Find 9u + 2v. With this question, do I multiply the u's by 9 and the v's by 2 then add like parts?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

if it helps write the vectors as 9(3x -1y) + 2(-6x-6y) the resulting ax+by is the vector (a,b)

OpenStudy (amistre64):

i just stack them tho, that way the placement stays intact and i dont move things on accident

OpenStudy (amistre64):

9<3, -1> +2<-6, -6> 27, -9 -12, -12 --------- a , b

OpenStudy (anonymous):

So it turns into <27 + -12; -9 + -12> <15, -21>?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (amistre64):

you know the dot product equivalence?

OpenStudy (amistre64):

|u| |v| cos(a) = u.v

OpenStudy (anonymous):

So how would I set this up? <20, 12> cos a?

OpenStudy (amistre64):

you know adding is adding like parts dot product is multiplying like parts, and adding the results

OpenStudy (amistre64):

|u| |v| cos(a) = 20+12 what are the length if u and v?

OpenStudy (anonymous):

u is 20 and v is 12

OpenStudy (amistre64):

no, the lenghts if anything sqrt(u.u) and sqrt(v.v)

OpenStudy (amistre64):

u = <-5, -4> |u| = sqrt(u.u) = sqrt(25+16) v = <-4, -3> |v| = sqrt(v.v) = sqrt(16+9)

OpenStudy (amistre64):

so 5sqrt(41) cos(a) = 20+12 how do we find the value of a?

OpenStudy (anonymous):

Divide the other side by 5sqrt41

OpenStudy (anonymous):

and do inverse cosine?

OpenStudy (amistre64):

yep, and cos^(-1) both sides a = cos^(-1) (32/(5sqrt(41))) now you want degrees, so make sure you either convert it, or set your calc on deg mode

OpenStudy (anonymous):

Okay, I got 1.78.

OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24> What would I do here to begin solving?